OpenStudy (anonymous):

Will someone please just tell me which one it is?! Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = 3x2 - 1

OpenStudy (anonymous):

3x2 or 3x^2

OpenStudy (anonymous):

im pretty sure it's f(x_ = 3x^2 - 1 @optiquest

OpenStudy (acxbox22):

Test to determine if a function y=f(x) is even, odd or neither: Replace x with -x and compare the result to f(x). If f(-x) = f(x), the function is even. If f(-x) = - f(x), the function is odd. sorce:google

OpenStudy (anonymous):

do that unless you want to do a proof

OpenStudy (anonymous):

@sammixboo @familyguymath @Hoslos @dtan5457 @Jhannybean @BassCatcher15 @BellaWolfe @iambatman @acxbox22 @freckles @familyguymath @superdavesuper @rational @whydoihavetosignup1 @Quxxn @Lady.Liv1776 @YanaSidlinskiy @IvantheBear @Openstudylover100 @GretaKnows @Gemini_Nation_ *Please someone just give me an answer?!?!?!*

OpenStudy (yanasidlinskiy):

Direct answers are against the Code of Conduct. You might would like to check out @e.mccormick 's post about direct answers. http://openstudy.com/users/e.mccormick#/updates/5511c987e4b08daf89890ff8

OpenStudy (anonymous):

I mean I'd enjoy an explanation but I usually take what I know and if I have the answer I can figure out how to do it..it's a bit odd I know but okay thank you

OpenStudy (yanasidlinskiy):

An explanation has already been posted by acboxx.